OpenStudy (anonymous):

6 – 3(2x – 4) = 2(1 + x) A. –2 B. 0 C. –1 D. 2

OpenStudy (anonymous):

Distribute 3 into the parenthesis. Can you do that?

OpenStudy (anonymous):

it will be 6-6X-12=2+2x ??

rishavraj (rishavraj):

@rocky1718 its supposed to be +12 .....:)

OpenStudy (anonymous):

It is? @rishavraj

OpenStudy (anonymous):

why ?

rishavraj (rishavraj):

6 – 3(2x – 4) = 6 - 6x + 12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

He's right. :P

OpenStudy (imtiaz7):

-3(2x – 4) = 2(1 + x) -6x+12=2+2x -6x-2x=2-12 -8x=-10 x=-10/-8 x=5/4 or x=1.25

OpenStudy (anonymous):

@Imtiaz7 Please do not give direct answers. They are against OpenStudy's Code of Conduct. If you keep giving them, you might get suspended. Please refrain from giving direct answers in the future. Thanks. :) If you do not know the rules, then you can read them here: http://openstudy.com/code-of-conduct

rishavraj (rishavraj):

@AnswerMyQuestions wht u meant..???o_O

OpenStudy (anonymous):

why is +12 ??

OpenStudy (imtiaz7):

bc - multiplied by - sign it comes +ve

OpenStudy (anonymous):

?

rishavraj (rishavraj):

https://www.mathsisfun.com/multiplying-negatives.html @rocky1718

OpenStudy (anonymous):

@rocky1718 So far, we have 6 - 6x + 12 = 2 + 2x, so can you figure it out?

OpenStudy (anonymous):

i subtract 12 and 6 ??

rishavraj (rishavraj):

@rocky1718 why subtract?? both r having positive sign ...they r supposed to be added...and did u understand why its +12????

rishavraj (rishavraj):

done?????

OpenStudy (anonymous):

alright i got it

OpenStudy (anonymous):

yeah